It's urgent please. Can anyone give the appropriate answer? This will carry 6 marks. Please help me.

1. (a) What is Supply law? Write down the factors that shift the supply curve. 

(b) What are the factors that determine elasticity of any good? 

(c) Suppose that the market demand for oranges is given by Q = 1000 – 250P and the market supply of oranges is given by Q = 150P, where P is the price of oranges and Q is the number of oranges. 

i) What are the equilibrium price and quantity for oranges? Illustrate your answer graphically. 

(ii) What will happen to equilibrium price and quantity of oranges if the price of income of our consumer increases? 

(iii) What will happen to equilibrium price and quantity of oranges if seeds of oranges become very expensive?
Answer #1

1 [ Demand p = 210 - Q Supply, p=20 Im equilibrium, Demand = Supply 210 - Q = 28 2 Q = 70 P = 140 ᏁᎢ , ᏗᎩ , , Ꭻ , ᎰᎢ ᎬᏔ ) 11-- This right ward shift of demand come will increase both price and quantity of tie in the equilibrium. Price will change be

answered by: ANURANJAN SARSAM